OpenStudy (anonymous):

What is the value of the expression when a = –4 and b = 5? (a+2b)^3 + 3a^2 A. –952 B. 168 C. –348 D. 264 @lacrosseplayer22 @Dscdago

OpenStudy (anonymous):

OO. Gurl. did you substitute the numbers in for the variables?

OpenStudy (anonymous):

no .-. I just did a huge problem kinda like this but im stuck all over again

OpenStudy (anonymous):

Substitute the numbers. So it looks like (-4+2(5))^3+3(-4)^2

OpenStudy (anonymous):

simply plug in a--4 b=5 you will get result i.e.264

OpenStudy (anonymous):

see it below how to do...

OpenStudy (anonymous):

so D?

OpenStudy (anonymous):

\[(-4+2(5))^3+3(-2)^2\]

OpenStudy (anonymous):

Correct. D it is.

OpenStudy (anonymous):

(-4+5*2)^3+3(-4)^2=(-4+10)^3+3(16)=6^3+48=216+48=264